# Positive mass and isoperimetry for continuous metrics with nonnegative scalar curvature

### v2

@@ 1.1. Main results
The classical positive mass theorem (PMT) in $3$ dimensions, originally due to Schoen–Yau [SchoenYauPMT], asserts that the ADM mass of a $3$-dimensional smooth complete asymptotically flat manifold with nonnegative scalar curvature is nonnegative. This seminal result has been later rediscovered, generalized, improved and exploited in a number of ways. Without any attempt to be complete, we refer to [WittenPMT, SYDefinitivePMT, AMO21, bray-karazas-khuri-stern] for different proofs and extensions to higher dimensions, and to [HuiskenIlmanen, BrayPenrose] for its main refinement, i.e., the Riemannian Penrose inequality.
@@
By a result of Gromov [GromovScal] (see also [BamlerGromov] for a different proof using Ricci flow) a smooth Riemannian manifold has nonnegative scalar curvature if and only if $R_{g}\geq 0$ holds in the approximate sense on $M$.
For a Borel set $E\subset M$, we let $|E|,P(E)$ denote the volume and the perimeter of $E$ respectively, see Section 2.2 for precise definitions. Let us recall that the perimeter of $E$ never charges the boundary of $M$, if any. The following quantity is a localized version of the isoperimetric mass introduced by Huisken [Huisken].
[…]
@@
$\begin{split}\mathfrak{m}_{\mathrm{iso}}\vcentcolon=\sup\bigg{\{}\limsup_{j\to+\infty}\mathfrak{m}_{\mathrm{QL}}(E_{j})\,\,\ :\ \,\,&E_{j}\subset M,P(E_{j})<+\infty\,\,\forall\,j,\,P(E_{j})\xrightarrow[]{}+\infty\bigg{\}}.\end{split}$ (1.2)
For asymptotically flat $3$-manifolds with nonnegative scalar curvature, the isoperimetric mass coincides with the ADM mass whenever the boundary is minimal [JaureguiLee, Theorem 3], [ChodoshEichmairShiYu21, Theorem C.2] (see also [BFMIsoperimetricRiemannianPenrose, Theorem 1.4] for a proof in the sharp asymptotic regime).
Huisken has conjectured [Huisken2] that a weak (isoperimetric) PMT should hold true for continuous Riemannian metrics if one interprets the notion of nonnegative scalar curvature (shortly, $R_{g}\geq 0$) in an appropriate weak sense. On the other hand in [JaureguiLeeUnger24IsoperimetricMass], after the first version of this preprint appeared, the authors proved that the property of being $C^{0}_{\rm loc}$-asymptotic to $\mathbb{R}^{n}$ (see Section 1.1) is enough to imply $\mathfrak{m}_{\mathrm{iso}}\geq 0$ in every dimension, without further imposing conditions on the curvature. The following is the precise meaning of $C^{0}_{\rm loc}$-asymptoticity.
@@
Let $K\subset M$ be a compact set (possibly empty) of a $C^{0}$-Riemannian manifold $(M,g)$. We say that an unbounded connected component $\mathcal{E}$ of $M\setminus K$ is $C^{0}_{\mathrm{loc}}$-asymptotic to a $C^{0}$-Riemannian manifold $(N,h)$ if the following holds. For every diverging sequence $\mathcal{E}\ni p_{i}\to+\infty$ there exists a point $o\in N$ such that $(M,g_{i},p_{i})\to(N,h,o)$ in the $C^{0}$-sense, see Section 2.1.
@@
Let $(M,g)$ be a complete 3-dimensional $C^{0}$-Riemannian manifold without boundary. Let $K\subset M$ be a compact set, and let $\mathcal{E}$ be an unbounded connected component $M\setminus K$. Assume that $\mathcal{E}$ is $C^{0}_{\mathrm{loc}}$-asymptotic to $\mathbb{R}^{3}$, see Section 1.1, and that $R_{g}\geq 0$ in the approximate sense on $\mathcal{E}\setminus K^{\prime}$, see Section 1.1, where $K^{\prime}\subset M$ is a compact set. Then for any ${C}>0$, there exists $E\subset\mathcal{E}\setminus K^{\prime}$ such that
@@
Of course (1.3) directly implies $\mathfrak{m}_{\mathrm{iso}}\geq 0$. More importantly, while the latter has been recently shown to hold in asymptotically flat manifolds regardless of any curvature condition [JaureguiLeeUnger24IsoperimetricMass], it is not difficult to construct asymptotically flat $3$-manifolds $(M,g)$ such that $\mathfrak{m}_{\rm QL}(E)<0$ for any bounded $E\subset M$. A general class of such manifolds is given by Cartan–Hadamard asymptotically flat $3$-manifolds with strictly negative sectional curvatures. In fact, any bounded domain sitting in such a space is known to satisfy the sharp Euclidean isoperimetric inequality [kleiner-isoperimetric] with strict inequality.
The same strategy used to prove Theorem 1.1 also yields a quasi-local isoperimetric PMT for small sets, without any asymptotic condition.
@@
As already suggested by Huisken [Huisken2], the nonnegativity of the isoperimetric mass should be compared with the following asymptotic formula, that can be directly deduced from [gray, Theorem 3.1]: there exists a positive constant $c$ such that for any smooth $3$-dimensional Riemannian manifold $(M,g)$, and any $o\in M$, it holds
@@
The latter (1.4) implies that, if ${R}_{g}(o)>0$, then $\mathfrak{m}_{\mathrm{QL}}(B_{r}(o))>0$ for sufficiently small $r$. However, this asymptotic reasoning does not provide a method to construct sets with nonnegative quasi-local mass in a smooth manifold with scalar curvature $\geq 0$, let alone sets for which the condition $\mathfrak{m}_{\mathrm{QL}}\geq 0$ remains stable under $C^{0}$-limits. This is what Theorem 1.2 instead provides in dimension $3$. Taking into account (1.4) we wonder whether the existence of arbitrarily small sets with nonnegative quasi-local mass could in fact characterize nonnegative scalar curvature. If this were the case, it would yield a new characterization of nonnegative scalar curvature that by Theorem 1.2 would be stable with respect to $C^{0}$-limits of smooth manifolds.
Both Theorem 1.1 and Theorem 1.2 are new even in the smooth setting, where they were known only under additional asymptotic assumptions, ensuring the existence of a global, proper weak inverse mean curvature flow [HuiskenIlmanen]. For general conditions under which such a flow exists, see [KaiXu]. In this case, the sets with nonnegative quasi-local mass are given by domains evolving through the inverse mean curvature flow, as a consequence of a reverse isoperimetric inequality one could infer from the work of Shi [Shi]. Similarly, our proof will rely on both a new localized version of the weak inverse mean curvature flow, thus allowing to drop the asymptotic assumptions, and on quantitative estimates that are stable for $C^{0}$-limits, under a lower bound on the scalar curvature.
The nonnegativity of the isoperimetric quasi-local mass for both small and large sets also has implications on the existence of isoperimetric regions. We show that in the $C^{0}$-setting with nonnegative scalar curvature, isoperimetric sets exist for sequences of both small and large volumes. The result below should be compared with [CarlottoChodoshEichmair, Proposition K.1].
@@
Let $(M,g)$ be a complete 3-dimensional $C^{0}$-Riemannian manifold without boundary, and assume that $R_{g}\geq 0$ in the approximate sense on $M\setminus\mathcal{C}$, where $\mathcal{C}$ is a compact set, see Section 1.1. Assume in addition that $M$ is $C^{0}_{\mathrm{loc}}$-asymptotic to $\mathbb{R}^{3}$, see Section 1.1.
@@
Very much like Theorem 1.1, also Theorem 1.3 strongly weakens the asymptotic assumptions for the existence of isoperimetric sets even in the smooth setting. The literature on the subject is very vast, in particular in relation to the study of canonical foliations of stable CMC compact hypersurfaces: we refer to the seminal [HuiskenYauFoliations] and to [CarlottoChodoshEichmair, eichmair-metzger_jdg, EichmairMetzger, nerz, ChodoshEichmairShiYu21, yu, eichmair-koerber, sinestrari-tenan], as well as to the references therein, for a fairly complete picture.
In the setting of Theorem 1.3 it is an interesting open problem to analyze existence of isoperimetric sets for any volume, uniqueness or foliation properties of such isoperimetric sets, as in [nerz, ChodoshEichmairShiYu21, yu] (see also the survey [BFSurvey]). Under the additional assumption that the isoperimetric profile $I$ is strictly increasing, we can actually prove existence of isoperimetric sets for every volume, see Section 4.2.
@@ 1.2. Strategy
We briefly discuss the strategy of the proof of Theorem 1.1, and Theorem 1.3, referring the reader to Section 3, and Section 4 for the details.
The starting point in the proof of Theorem 1.1 and Theorem 1.2 is the following consequence of a result due to Shi [Shi], after important insights by Brendle–Chodosh [brendle-chodosh]: in a smooth $3$-dimensional asymptotically flat manifold with nonnegative scalar curvature the level sets of the weak inverse mean curvature flow (shortly, IMCF, see Section 3) issuing from a point satisfy a reverse Euclidean isoperimetric inequality with the sharp constant, as long as their boundaries are connected. Hence the isoperimetric deficit in (1.1) is nonnegative when computed on such sets, and this directly provides examples of arbitrarily large regions with nonnegative quasi-local mass.
We will push this heuristic to show that on the approximating manifolds $(M,g_{j})$ one can define a well-behaved local weak IMCF $w_{j}$ on punctured balls $B_{j}$. This is done by taking scaled limits of the logarithms of $p$-Green functions on such balls, for $p\to 1^{+}$, as pioneered by R. Moser [MoserIMCF]. In fact, we build on the sharp gradient estimate obtained in [kotschwar_localgradientestimatesharmonic_2009], and on the Harnack inequality with explicit constants in [salvatori-rigoli-vignati] to get such functions $w_{j}$ with bounds from below only in terms of constants that stay bounded in $j$ when the $B_{j}$’s are $C^{0}$-close enough to a Euclidean ball. In general, these estimates will be in force for balls $B_{j}$ that are small enough, while under additional $C^{0}_{\rm loc}$-asymptotics, they can be obtained for arbitrarily large balls $B_{R_{j}}(p_{j})$ with $R_{j}\to\infty$ and $p_{j}\to\infty$. This is in fact the content of the quantitative existence result for local IMCFs Theorem 3.1. The latter finally allows to pass to the limit the level sets of these IMCFs, obtaining the sets satisfying the (sharp) reverse Euclidean isoperimetric inequality (or equivalently with nonnegative quasi-local mass) claimed in Theorem 1.1 and Theorem 1.2, respectively.
The proof of the existence of isoperimetric sets in Theorem 1.3 is not constructive, and it is based on a contradiction argument. One notices that if after a certain volume threshold isoperimetric sets do not exist, then the isoperimetric profile is strictly increasing for large volumes: this is a consequence of a generalized existence theorem for the isoperimetric problem, see Theorem 2.2. In addition, arguing as in the previous paragraph, we construct sets that satisfy the reverse sharp Euclidean isoperimetric inequality with arbitrarily large volumes and perimeters, and that avoid any fixed compact set. This is enough to show, again using Theorem 2.2 and that the isoperimetric profile is increasing, that isoperimetric sets with arbitrarily large volumes must exist, thus resulting in a contradiction. The analogous statement for small volumes follows from an easier instance of this reasoning.
@@ 1.3. Comments and comparison with related literature
1.3.1. Other notions of weak scalar curvature bounds, and relations with the works [PaulaGAFA, PaulaADM]
Several notions of scalar curvature lower bounds for smooth manifolds endowed with $C^{0}$-Riemannian metrics have been proposed in the recent years. In [GromovScal] Gromov has proposed a definition based on nonexistence of suitable small polyhedra on the manifold, see the recent works [ChaoLiPolyedron, BrendleInv] motivated by this study; a definition based on regularization through Ricci flow has been suggested by Burkhardt-Guim in [PaulaGAFA]; assuming the nonnegativity of the right hand side of (1.4) as a replacement for nonnegative scalar curvature has been hinted at by Huisken [Huisken2]. Let us compare now nonnegative scalar curvature in the approximate sense with the notion in [PaulaGAFA, PaulaADM].
Assume that a complete $C^{0}$-Riemannian manifold $(M,g)$ is globally $C^{0}$-asymptotic to $\mathbb{R}^{3}$ outside a compact set $K\subset M$, i.e. $M\setminus K=\mathbb{R}^{n}\setminus B$ for some ball $B$ and $g=\delta+o(1)$ at infinity, with $\delta$ the flat metric. Hence, for $\beta\in(0,1/2)$, we claim that if $R_{g}\geq 0$ in the $\beta$-weak sense [PaulaADM, Definition 2.3] on $M\setminus K$, then $R_{g}\geq 0$ in the approximate sense on $M\setminus K$, up to possibly enlarging $K$. Indeed, this is due to the fact that under the $C^{0}$-asymptotic hypothesis at infinity one can first define a Ricci-deTurck flow $(M,g_{t})$ starting from $g$ which $C^{0}$-converges to $(M,g)$ locally uniformly by using [SimonC0, Theorem 1.1]. Then, since $R_{g}\geq 0$ in the $\beta$-weak sense on $M\setminus K$, [PaulaADM, Lemma 5.1] implies that, up to possibly enlarging $K$, $R_{g_{t}}\geq-o(1)$ on $M\setminus K$ as $t\to 0$, which in turn implies that $R_{g}\geq 0$ in the approximate sense on $M\setminus K$ by definition. In the case $M$ is a compact manifold, the previous reasoning has been explicitly recorded in [PaulaGAFA, Corollary 1.6]. Related to this, it would be interesting to understand whether a non-compact $C^{0}$-manifold $(M,g)$ that has $R_{g}\geq 0$ in the approximate sense on $M$ admits a smooth metric $\tilde{g}$ with $R_{\tilde{g}}\geq 0$. The compact case has been settled in [PaulaGAFA, Corollary 1.8]. Taking into account the discussion above, this seems likely to hold when $(M,g)$ is, in addition, $C^{0}$-asymptotic to $\mathbb{R}^{n}$ globally.
As a consequence of the above discussion, under the stronger assumption that $(M,g)$ is globally $C^{0}$-asymptotically flat one gets that our theorems 1.1, 1.2, and 1.3 hold under the assumption that $R_{g}\geq 0$ in the $\beta$-weak sense outside a compact set, for some $\beta\in(0,1/2)$ ([PaulaADM, Definition 2.3]).
Finally, we point out that as of today it is not known whether the definitions in [GromovScal] and [PaulaGAFA] are equivalent, compare with the discussion [PaulaGAFA, Page 1707-1708], nor any relation with Huisken’s notion in [Huisken2] has been investigated yet.

### v3

@@ 1.1. Main results
The classical positive mass theorem (PMT) in $3$ dimensions, originally due to Schoen–Yau [67], asserts that the ADM mass of a $3$-dimensional smooth complete asymptotically flat manifold with nonnegative scalar curvature is nonnegative. This seminal result has been later rediscovered, generalized, improved and exploited in a number of ways. Without any attempt to be complete, we refer to [72, 68, 1, 15] for different proofs and extensions to higher dimensions, and to [38, 16] for its main refinement, i.e., the Riemannian Penrose inequality.
@@
By a result of Gromov [33] (see also [8] for a different proof using Ricci flow) a smooth Riemannian manifold has nonnegative scalar curvature if and only if $R_{g}\geq 0$ holds in the approximate sense on $M$.
For a Borel set $E\subset M$, we let $|E|,P(E)$ denote the volume and the perimeter of $E$ respectively, see Section˜2.2 for precise definitions. Let us recall that the perimeter of $E$ never charges the boundary of $M$, if any. The following quantity is a localized version of the isoperimetric mass introduced by Huisken [40].
Definition 1.2 (Quasi-local isoperimetric mass).
@@
For asymptotically flat $3$-manifolds with nonnegative scalar curvature, the isoperimetric mass coincides with the ADM mass whenever the boundary is minimal [43, Theorem 3], [25, Theorem C.2] (see also [11, Theorem 1.4] for a proof in the sharp asymptotic regime).
Huisken has conjectured [41] that a weak (isoperimetric) PMT should hold true for continuous Riemannian metrics if one interprets the notion of nonnegative scalar curvature (shortly, $R_{g}\geq 0$) in an appropriate weak sense. On the other hand in [42], after the first version of this preprint appeared, the authors proved that the property of being $C^{0}_{\rm loc}$-asymptotic to $\mathbb{R}^{n}$ (see Section˜1.1) is enough to imply $\mathfrak{m}_{\mathrm{iso}}\geq 0$ in every dimension, without further imposing conditions on the curvature. The following is the precise meaning of $C^{0}_{\rm loc}$-asymptoticity.
Definition 1.3.
Let $K\subset M$ be a compact set (possibly empty) of a $C^{0}$-Riemannian manifold $(M,g)$. We say that an unbounded connected component $\mathcal{E}$ of $M\setminus K$ is $C^{0}_{\mathrm{loc}}$-asymptotic to a $C^{0}$-Riemannian manifold $(N,h)$ if the following holds. For every diverging sequence $\mathcal{E}\ni p_{i}\to+\infty$ there exists a point $o\in N$ such that $(M,g_{i},p_{i})\to(N,h,o)$ in the $C^{0}$-sense, see Section˜2.1.
@@
Let $(M,g)$ be a complete 3-dimensional $C^{0}$-Riemannian manifold without boundary. Let $K\subset M$ be a compact set, and let $\mathcal{E}$ be an unbounded connected component $M\setminus K$. Assume that $\mathcal{E}$ is $C^{0}_{\mathrm{loc}}$-asymptotic to $\mathbb{R}^{3}$, see Section˜1.1, and that $R_{g}\geq 0$ in the approximate sense on $\mathcal{E}\setminus K^{\prime}$, see Section˜1.1, where $K^{\prime}\subset M$ is a compact set. Then for any C>0{C}>0, there exists $E\subset\mathcal{E}\setminus K^{\prime}$ such that
@@
Of course (1.3) directly implies $\mathfrak{m}_{\mathrm{iso}}\geq 0$. More importantly, while the latter has been recently shown to hold in asymptotically flat manifolds regardless of any curvature condition [42], it is not difficult to construct asymptotically flat $3$-manifolds $(M,g)$ such that $\mathfrak{m}_{\rm QL}(E)<0$ for any bounded $E\subset M$. A general class of such manifolds is given by Cartan–Hadamard asymptotically flat $3$-manifolds with strictly negative sectional curvatures. In fact, any bounded domain sitting in such a space is known to satisfy the sharp Euclidean isoperimetric inequality [46] with strict inequality.
The same strategy used to prove Theorem˜1.4 also yields a quasi-local isoperimetric PMT for small sets, without any asymptotic condition.
@@
As already suggested by Huisken [41], the nonnegativity of the isoperimetric mass should be compared with the following asymptotic formula, that can be directly deduced from [31, Theorem 3.1]: there exists a positive constant $c$ such that for any smooth $3$-dimensional Riemannian manifold $(M,g)$, and any $o\in M$, it holds
@@
The latter (1.4) implies that, if ${R}_{g}(o)>0$, then $\mathfrak{m}_{\mathrm{QL}}(B_{r}(o))>0$ for sufficiently small $r$. However, this asymptotic reasoning does not provide a method to construct sets with nonnegative quasi-local mass in a smooth manifold with nonnegative scalar curvature, let alone sets for which the condition $\mathfrak{m}_{\mathrm{QL}}\geq 0$ remains stable under $C^{0}$-limits. This is what Theorem˜1.5 instead provides in dimension $3$. Taking into account (1.4) we wonder whether the existence of arbitrarily small sets with nonnegative quasi-local mass could in fact characterize nonnegative scalar curvature. If this were the case, it would yield a new characterization of nonnegative scalar curvature that by Theorem˜1.5 would be stable with respect to $C^{0}$-limits of smooth manifolds.
Both Theorem˜1.4 and Theorem˜1.5 are new even in the smooth setting, where they were known only under additional asymptotic assumptions, ensuring the existence of a global, proper weak inverse mean curvature flow [38]. For general conditions under which such a flow exists, see [73]. In this case, the sets with nonnegative quasi-local mass are given by domains evolving through the inverse mean curvature flow, as a consequence of a reverse isoperimetric inequality one could infer from the work of Shi [69]. Similarly, our proof will rely on both a new localized version of the weak inverse mean curvature flow, thus allowing to drop the asymptotic assumptions, and on quantitative estimates that are stable for $C^{0}$-limits, under a lower bound on the scalar curvature.
The nonnegativity of the isoperimetric quasi-local mass for both small and large sets also has implications on the existence of isoperimetric regions. We show that in the $C^{0}$-setting with nonnegative scalar curvature, isoperimetric sets exist for sequences of both small and large volumes. The result below should be compared with [23, Proposition K.1].
@@
Let $(M,g)$ be a complete 3-dimensional $C^{0}$-Riemannian manifold without boundary, and assume that $R_{g}\geq 0$ in the approximate sense on $M\setminus\mathcal{C}$, where $\mathcal{C}$ is a compact set, see Section˜1.1. Assume in addition that $M$ is $C^{0}_{\mathrm{loc}}$-asymptotic to $\mathbb{R}^{3}$, see Section˜1.1.
@@
Very much like Theorem˜1.4, also Theorem˜1.6 strongly weakens the asymptotic assumptions for the existence of isoperimetric sets even in the smooth setting. The literature on the subject is very vast, in particular in relation to the study of canonical foliations of stable CMC compact hypersurfaces: we refer to the seminal [39] and to [23, 27, 28, 60, 25, 75, 26, 71], as well as to the references therein, for a fairly complete picture.
In the setting of Theorem˜1.6 it is an interesting open problem to analyze existence of isoperimetric sets for any volume, uniqueness or foliation properties of such isoperimetric sets, as in [60, 25, 75] (see also the survey [12]). Under the additional assumption that the isoperimetric profile $I$ is strictly increasing, we can actually prove existence of isoperimetric sets for every volume, see Section˜4.2.
@@ 1.2. Strategy
We briefly discuss the strategy of the proof of Theorem˜1.4, and Theorem˜1.6, referring the reader to Section˜3, and Section˜4 for the details.
The starting point in the proof of Theorem˜1.4 and Theorem˜1.5 is the following consequence of a result due to Shi [69], after important insights by Brendle–Chodosh [17]: in a smooth $3$-dimensional asymptotically flat manifold with nonnegative scalar curvature the level sets of the weak inverse mean curvature flow (shortly, IMCF, see Section˜3) issuing from a point satisfy a reverse Euclidean isoperimetric inequality with the sharp constant, as long as their boundaries are connected. Hence the isoperimetric deficit in (1.1) is nonnegative when computed on such sets, and this directly provides examples of arbitrarily large regions with nonnegative quasi-local mass.
We will push this heuristic to show that on the approximating manifolds $(M,g_{j})$ one can define a well-behaved local weak IMCF $w_{j}$ on punctured balls $B_{j}$. This is done by taking scaled limits of the logarithms of $p$-Green functions on such balls, for $p\to 1^{+}$, as pioneered by R. Moser [56]. In fact, we build on the sharp gradient estimate obtained in [47], and on the Harnack inequality with explicit constants in [64] to get such functions $w_{j}$ with bounds from below only in terms of constants that stay bounded in $j$ when the $B_{j}$’s are $C^{0}$-close enough to a Euclidean ball. In general, these estimates will be in force for balls $B_{j}$ that are small enough, while under additional $C^{0}_{\rm loc}$-asymptotics, they can be obtained for arbitrarily large balls $B_{R_{j}}(p_{j})$ with $R_{j}\to\infty$ and $p_{j}\to\infty$. This is in fact the content of the quantitative existence result for local IMCFs Theorem˜3.6. The latter finally allows to pass to the limit the level sets of these IMCFs, obtaining the sets satisfying the (sharp) reverse Euclidean isoperimetric inequality (or equivalently with nonnegative quasi-local mass) claimed in Theorem˜1.4 and Theorem˜1.5, respectively.
The proof of the existence of isoperimetric sets in Theorem˜1.6 is not constructive, and it is based on a contradiction argument. One notices that if after a certain volume threshold isoperimetric sets do not exist, then the isoperimetric profile is strictly increasing for large volumes: this is a consequence of a generalized existence theorem for the isoperimetric problem, see Theorem˜2.16. In addition, arguing as in the previous paragraph, we construct sets that satisfy the reverse sharp Euclidean isoperimetric inequality with arbitrarily large volumes and perimeters, and that avoid any fixed compact set. This is enough to show, again using Theorem˜2.16 and that the isoperimetric profile is increasing, that isoperimetric sets with arbitrarily large volumes must exist, thus resulting in a contradiction. The analogous statement for small volumes follows from an easier instance of this reasoning.
@@ 1.3. Comments and comparison with related literature
1.3.1. Other notions of weak scalar curvature bounds, and relations with the works [19, 20]
Several notions of scalar curvature lower bounds for smooth manifolds endowed with $C^{0}$-Riemannian metrics have been proposed in the recent years. In [33] Gromov has proposed a definition based on nonexistence of suitable small polyhedra on the manifold, see the recent works [50, 18] motivated by this study; a definition based on regularization through Ricci flow has been suggested by Burkhardt-Guim in [19]; assuming the nonnegativity of the right hand side of (1.4) as a replacement for nonnegative scalar curvature has been hinted at by Huisken [41]. Let us compare now nonnegative scalar curvature in the approximate sense with the notion in [19, 20].
Assume that a complete $C^{0}$-Riemannian manifold $(M,g)$ is globally $C^{0}$-asymptotic to $\mathbb{R}^{3}$ outside a compact set $K\subset M$, i.e. $M\setminus K=\mathbb{R}^{n}\setminus B$ for some ball $B$ and $g=\delta+o(1)$ at infinity, with $\delta$ the flat metric. Hence, for $\beta\in(0,1/2)$, we claim that if $R_{g}\geq 0$ in the $\beta$-weak sense [20, Definition 2.3] on $M\setminus K$, then $R_{g}\geq 0$ in the approximate sense on $M\setminus K$, up to possibly enlarging $K$. Indeed, this is due to the fact that under the $C^{0}$-asymptotic hypothesis at infinity one can first define a Ricci-deTurck flow $(M,g_{t})$ starting from $g$ which $C^{0}$-converges to $(M,g)$ locally uniformly by using [70, Theorem 1.1]. Then, since $R_{g}\geq 0$ in the $\beta$-weak sense on $M\setminus K$, [20, Lemma 5.1] implies that, up to possibly enlarging $K$, $R_{g_{t}}\geq-o(1)$ on $M\setminus K$ as $t\to 0$, which in turn implies that $R_{g}\geq 0$ in the approximate sense on $M\setminus K$ by definition. In the case $M$ is a compact manifold, the previous reasoning has been explicitly recorded in [19, Corollary 1.6]. Related to this, it would be interesting to understand whether a non-compact $C^{0}$-manifold $(M,g)$ that has $R_{g}\geq 0$ in the approximate sense on $M$ admits a smooth metric $\tilde{g}$ with $R_{\tilde{g}}\geq 0$. The compact case has been settled in [19, Corollary 1.8]. Taking into account the discussion above, this seems likely to hold when $(M,g)$ is, in addition, $C^{0}$-asymptotic to $\mathbb{R}^{n}$ globally.
As a consequence of the above discussion, under the stronger assumption that $(M,g)$ is globally $C^{0}$-asymptotically flat one gets that our theorems 1.4, 1.5, and 1.6 hold under the assumption that $R_{g}\geq 0$ in the $\beta$-weak sense outside a compact set, for some $\beta\in(0,1/2)$ ([20, Definition 2.3]).
Finally, we point out that as of today it is not known whether the definitions in [33] and [19] are equivalent, compare with the discussion [19, Page 1707-1708], nor any relation with Huisken’s notion in [41] has been investigated yet.